OpenStudy (anonymous):

Medal for Help ASAP :)What is an Expression for the sale price of a bracelet that has been discounted 60% from it's sticker price ? Evaluate the expression for the original sticker price $90? A. s= p - 0.6p;$36 B. s= p - 0.4p; $54 **** My answer C. s= p - 60p; $5,490 D. s= p - 0.6p; $90

OpenStudy (anonymous):

@Clalgee

OpenStudy (anonymous):

@kropot72

OpenStudy (clalgee):

Alright :) So for this question, my answer would be S = p - 0.6p -- p = 90 because you need to get 60% from the original price of $90.

OpenStudy (clalgee):

'S = p - 0.6p -- p = 90' or D in this case :)

OpenStudy (anonymous):

Ok I get it Ty I have one more can u help :)

OpenStudy (clalgee):

Yeah! Sure :)

OpenStudy (anonymous):

Ok going to post on the MATH Wall and tag u

OpenStudy (clalgee):

Alright! Sounds good :)

OpenStudy (anonymous):

@Clalgee I luv u so much bae. Be muy bar